OpenStudy (anonymous):

PLEASE HELP: Suzy grabs four marbles from a bag containing four white marbles and five blue marbles. Find the probability that exactly three of the marbles are blue.

OpenStudy (kropot72):

\[P(4\ blue)=\frac{\left(\begin{matrix}5 \\ 3\end{matrix}\right)\left(\begin{matrix}4 \\ 1\end{matrix}\right)}{\left(\begin{matrix}9 \\ 4\end{matrix}\right)}\]

OpenStudy (anonymous):

awesome! thank you, i have been struggling forever trying to figure out how to do this

OpenStudy (anonymous):

in situations how do you know when to use nCr or nPr

OpenStudy (kropot72):

np. As you can see, the hypergeometric probability distribution applies. Regarding you question, nCr deals with combinations where the order of selection does not matter. nPr deals with permutations where the order of selection is important. When calculating the hypergeometric probability, combinations are used.

OpenStudy (anonymous):

oh okay now i remember that! could i have you help me with another one really quick too?

OpenStudy (kropot72):

So the calculation for your question can be written as: \[P(3\ blue)=\frac{5C3\times4C1}{9C4}\]

OpenStudy (anonymous):

okay yeah thats how i typed it in the calculator

OpenStudy (kropot72):

Sorry, I have to log out now. Please post as a new Question :)

OpenStudy (anonymous):

okay thanks anyways! :)